An ideal Diesel cycle has a compression ratio of 17 and a cut-off ratio of 2. At the beginning of the compression process P = 100 kPa and T = 27◦C. Find the temperature, pressure and specific volume at states 1, 2, 3, and 4. Find the heat supplied, heat rejected, work done by the system, work done on the system, net work done, and thermal efficiency. Calculate the thermal efficiency of a Carnot cycle operating between the same temperature limits.
Answer #1

given P = 100kPa T= 27C = 300k AD/ Compression Ratio (0)=17 (s) Cut off Ratioe: 2 sp. heat Ratio 1) = 1.4 - > V-> at 1, by Id273. Process P= P2 = 5279.9 MPa] given cut off ratio €)= dg: 2 dz=24 00506 193 = 0,01 mes YAYAYAYAYAYANAN Applying Ideal gasMeat addition Dz.3 = MC T3-T2) . = 1X 7005 (1865.23 – 931675) P2 = 998.14 med 938:14 KS heat rejection - - = Jesu.1 = 0 C (-)work done by the system (w.) W = Warz + Way Pad. De) 4 RCT3-T4). = 5279.97(701-OSOL) + 287 (1865-23-991.5) 104-1 1036515|pgThermal efficiency (1) s Whet Gsupplied (923) . 938.14 5. = 8.631 11 - 63.). C3... 1 Thermal efficieny of Carnot engine wassi